# InterDyn: Controllable Interactive Dynamics with Video Diffusion Models

### v1

@@ 4.2 Metrics
Image quality. To assess per-frame image quality we report the Structural Similarity Index Measure (SSIM[76], Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS[100].
Spatio-temporal similarity. To assess the spatio-temporal perceptual similarity between the ground truth and the generated video distributions, we employ the Fréchet Video Distance (FVD[70], as implemented in [65].
Motion Fidelity. Since we do not explicitly control object dynamics, pixel alignment between the object in the ground truth and the generated video is only guaranteed in the starting frame. In light of this, we assess the quality of object motion by adapting the motion fidelity metric proposed by Yatim et al. [90], which measures the similarity between point-tracking trajectories.
Using the object bounding box from the Something-Else dataset as an input prompt to SAM2 [61], we obtain a mask of the object in the starting frame of the video. Using the mask, we sample 100 points on the object and track these throughout both the ground truth and generated video using CoTracker3 [38]. Given the resulting two sets of tracklets $\mathcal{T}=\{\tau_{1},\dots,\tau_{n}\},\tilde{\mathcal{T}}=\{\tilde{\tau}_{1},\dots,\tilde{\tau}_{m}\}$ the motion fidelity metric measures the correlation their and is defined as:
@@
where the correlation between two tracklets $\textbf{corr}(\tau,\tilde{\tau})$ [47] is defined as:
$\textbf{corr}(\tau,\tilde{\tau})\frac{1}{F}\sum_{k=1}^{F}\frac{{v_{k}^{x}\cdot\tilde{v}_{k}^{x}+v_{k}^{y}\cdot\tilde{v}_{k}^{y}}}{{\sqrt{(v_{k}^{x})^{2}+(v_{k}^{y})^{2}}\cdot\sqrt{(\tilde{v}_{k}^{x})^{2}+(\tilde{v}_{k}^{y})^{2}}}}\text{.}$ (2)
Here $(v_{k}^{x},v_{k}^{y}),(\tilde{v}_{k}^{x},\tilde{v}_{k}^{y})$ are the $k^{th}$ frame displacement of tracklets $\tau,\tilde{\tau}$ respectively. If there are less than 100 points to query on the object due to it being too small, we do not consider the video for the motion fidelity metric.
@@ 4.3 Probing dynamics with object collision events
We investigate the ability of SVD to generate interactive dynamics using the synthetic CLEVRER dataset [94], which includes 20,000 videos of colliding objects with annotated segmentation masks, 3D locations, and collision events. We sample videos where stationary objects interact with moving objects entering the scene. Segmentation masks are used to construct control signals for moving objects, and differently colored masks help the model distinguish unique objects. Frames are randomly cropped and scaled to $320\times 448$, and starting frames are sampled before collisions to maximize exposure to dynamics. Our primary goal is to assess whether InterDyn can generate future frames that are physically plausible in response to introduced interactions. Specifically, we examine whether the model can produce meaningful object movements for uncontrolled objects in the scene, given the movement of objects entering the scene—for instance, determining how an object would move if struck by another.
Beyond direct interactions, we focus on two more challenging physical scenarios: Counterfactual Dynamics: We investigate whether the model can predict plausible motion trajectories involving multiple interactions at different time steps. This involves generating realistic motions under counterfactual scenarios, where different interaction configurations lead to different causal effects—much like a physics simulator. This tests how multiple interactions influence a single object. Force Propagation: We examine whether the model can generate subsequent object dynamics through force propagation. This means assessing whether objects beyond our direct control interact with each other in physically plausible ways. Here, we are interested in how one interaction influences multiple objects. We conduct both experiment settings on the CLEVRER test set.
As illustrated in the “Future 1” row in Fig. 3(a), the gray cylinder (controlled object 1) collides with the stationary red sphere, causing it to move. Along its new trajectory, the red sphere is then struck by the brown cylinder (controlled object 2), altering its path once again. Compared to the top row, which shows the ground truth physics-simulated renderings, InterDyn synthesizes dynamics that are physically plausible without explicit knowledge of the objects’ mass or stiffness. For the counterfactual experiment shown in the “Future 2” row, we remove the brown cylinder and observe how dynamics change after this. Without this second interaction, the red sphere continues along its original trajectory, as indicated by the red point-tracking path—consistent with physical expectations. Importantly, there is no control signal for the red sphere throughout the sequence; its motion is entirely generated by InterDyn.
@@
InterDyn can generate force propagation dynamics, where uncontrolled objects interact with each other, as illustrated in Fig. 3(b). The sole driving force is the red cylinder at the top left. This controlled object moves towards and collides with the uncontrolled blue cylinder, setting it into motion. The blue cylinder then collides with the uncontrolled purple cylinder, which subsequently impacts the uncontrolled purple cube on the far right. The point-tracking trajectories display the position changes of all three objects in a manner consistent with physical laws. This experiment suggests that InterDyn possesses an implicit understanding of physical interactions, enabling it to generate plausible dynamic sequences. We provide directions to (more) results in video format in Appendix A.
While the CLEVRER experiments highlight the fundamental capabilities of InterDyn in generating interactive dynamics, we are also interested in evaluating its performance in complex, real-world scenarios—the environments for which the video model was originally trained.
@@ 4.4 Generating Human-Object Interactions
We find such scenarios in the Something-Something-v2 dataset [20]. Originally proposed for human action recognition and video understanding, this dataset provides 220,847 videos of humans performing basic actions with everyday objects. It contains actions like “pushing [something] from left to right”, “squeezing [something]” and “lifting [something] with [something] on it”. The Something-Else dataset [55] provides bounding box annotations for 180,049 of these videos, which we use to extract hand masks with SAM2. This dataset allows us to train InterDyn at a larger scale and to compare with our closest competitor CosHand [66]. We train one version at the same resolution as CosHand, $256\times 256$, and a second version at $256\times 384$, which aligns better with the video resolution SVD has been trained on. We provide directions to (more) results in video format in Appendix A.
[…]
4.4.1 Baselines
We compare InterDyn with two CosHand variants: a frame-by-frame approach and an auto-regressive approach. The results can be found in Tab. 1. In the first approach, we predict each frame $\hat{x}_{t+1}$ in the sequence from the input frame and its corresponding mask:
$\hat{x}_{t+1}=\text{\mbox{CosHand}}(x_{0},h_{0},h_{t+1})\text{,}\quad\forall t\in[1,13],$ (3)
where $h_{0},h_{t}$ denote a mask, and $\hat{x}_{0}$ is the initial frame. In the second approach, we use CosHand to generate videos auto-regressively in the following manner:
$\hat{x}_{t+1}=\text{\mbox{CosHand}}(\hat{x}_{t},h_{t},h_{t+1})\text{.}\quad\forall t\in[1,13],$ (4)
where $\hat{x}_{t}$ is the generated frame from the previous time step. To avoid error propagation from the VAE decoding-encoding process, $\hat{x}_{t+1}$ remains in latent space.
@@ 4.4.1 Baselines
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ Motion Fidelity $\uparrow$
CosHand 14-frame 0.615 16.87 0.313 91.18 0.432
CosHand Auto-regressive 0.531 14.92 0.408 90.30 0.570
Ours (256$\times$256) 0.664 18.60 0.260 19.27 0.633
Ours (256$\times$384) 0.680 19.04 0.252 22.22 0.641
Table 1: Comparisons on the Something-Something-v2 dataset (video data). We compare against two video extensions of CosHand: (1) by generating each frame from the initial image (CosHand 14-frame), (2) by auto-regressively generating the frames (CosHand Auto-regressive). All videos are generated with 14 frames. We provide two resolutions for InterDyn: 256$\times$256 (same resolution as CosHand) and 256$\times$384 (more compatible with SVD). We evaluate in terms of SSIM, PSNR, LPIPS, FVD, and a tracking-based metric (Motion Fidelity[90].
Analysis
@@
We observe that CosHand in the frame-by-frame setting achieves high image quality, as indicated by the SSIM, PSNR, and LPIPS metrics, but struggles with temporal coherence and motion fidelity. Qualitatively, we notice that the object location may appear inconsistently across frames due to the lack of temporal context. In contrast, the auto-regressive version improves object motion fidelity over the frame-by-frame approach but suffers from lower frame-wise image quality due to error propagation. Importantly, it fails in scenarios requiring accurate post-interaction dynamics, such as when objects continue moving after being released from direct hand contact as shown in Fig. 4.
Our method, InterDyn (256$\times$384), achieves the best overall performance, surpassing CosHand in image quality, spatio-temporal dynamics, and motion fidelity. It generates realistic post-interaction dynamics—such as rolling or sliding objects—and effectively handles noisy control signals. InterDyn (256$\times$384) aligns better with the frozen U-Net prior, outperforming the InterDyn (256$\times$256) variant.
Different from its two-states-based competitors, InterDyn interprets the control signals at the sequential motion level, rather than simply pixel-aligning with the provided hand masks. For instance, when SAM2 produces noisy mask sequences due to significant motion blur (as shown in Fig. 6), leading to incomplete or coarse hand shapes, InterDyn remains robust during inference. By leveraging its temporal prior and temporal control branch to handle noisy outliers, InterDyn can translate coarse and noisy masks into detailed hands –including individual fingers–, holistically reasoning about the temporal control signal. This sets it apart from methods that rely on precise input mask fidelity. We have conducted experiments on different variants for the control signal, see Appendix B.
[…]
Static comparison.
@@
Table 2: Comparisons on the Something-Something-v2 dataset (static data). We compare against CosHand and other static baselines on the task of generating a single future frame. The top five rows were taken from [66]. Since CosHand does not provide the exact validation split, we also evaluate their method on our own validation set. For a fair comparison, we train two methods: one with the same resolution as in CosHand (256$\times$256), and another one on the native resolution of SVD (256$\times$384).
For completeness, we also compare against CosHand [66] on a per-frame basis. For this, we extract the per-frame metrics for the second frame from InterDyn’s and CosHand’s video generation, and compare against the baselines evaluated in [66]. Quantitative results are shown in Tab. 2.
4.4.2 Diverse physical generation with InterDyn
We present several examples of the diverse interactive dynamics generated by InterDyn in Figure 5. Row 1 shows how InterDyn generates the articulated motion of an object. Row 2 showcases pouring water into a glass; note how the water level increases over time. Row 3 demonstrates an object being dropped, moving out of frame when falling, and rolling back in frame once hitting the floor, featuring realistic motion blur synthesis. Rows 4 and 5 illustrate how InterDyn handles squeezing interactions—the rubber and the spring are compressed and restored accordingly. Row 6 demonstrates an understanding of physical size and distance to the camera, as the phone moves closer to the viewer. These results highlight the complexity that InterDyn is capable of handling, implying its generalization ability and future potential as an implicit yet generalized physical simulator and renderer.
5 Conclusion
We introduced InterDyn, a framework that generates videos of interactive dynamics by leveraging large video foundation models as implicit physics simulators. By incorporating an interactive control mechanism, InterDyn produces plausible, temporally consistent videos of object interactions—including complex human-object interactions—while generalizing to unseen objects. Our evaluations demonstrate that InterDyn effectively captures continuous motion and subsequent dynamics, outperforming baselines that focus on single future states. This work highlights the potential of using video generative models for physics simulation without explicit reconstruction, opening new avenues for future research.
Conflict of interest disclosure: MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon and Meshcapade GmbH. While MJB is a co-founder and Chief Scientist at Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Supplementary Material
Appendix A Additional Results
We provide more qualitative results in Fig. 9 and as videos at: https://interdyn.is.tue.mpg.de/.
Appendix B Control Signal
In this work, we use binary masks of a hand as the control signal. However, other types of signals such as skeletons or meshes could provide richer controllability, since they inherently encode pseudo-3D information and capture correspondences across frames. Nevertheless, collecting large datasets with these types of annotations is significantly more challenging compared to binary masks, for which high-performance image segmentation models like SAM2 are readily available.
To test the impact of our choice, we ran an experiment on the DexYCB dataset [10]. Originally proposed for 3D pose estimation, DexYCB provides 8000 videos of hands grasping an object with accompanying ground-truth hand and object poses [10]. To represent 3D hand pose, DexYCB uses the parametric human hand model MANO [63], which we render as a binary mask, as joints in a similar style as OpenPose [8], and as a colored mesh using the colormap from [57], see Fig. 7.
We noticed that when generating hand masks with SAM2 on the Something-Something-v2 dataset, the resulting masks have the object’s contour cut out when held in the hand. This inadvertently provides InterDyn with a small amount of information about the shape of the object that is being manipulated (a limitation shared with CosHand [66]). For this experiment, to ensure a fair comparison between control signals, we train separate versions of InterDyn that explicitly incorporate or leave out occlusions of all objects in the scene, see Fig. 7.
[…]
Occl. SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ Motion Fidel. $\uparrow$
Mask 0.829 24.08 0.123 39.94 0.666
Joints 0.827 24.00 0.124 40.02 0.673
Mesh 0.828 24.14 0.122 41.99 0.663
Mask 0.829 24.15 0.122 37.64 0.675
Joints 0.827 24.05 0.124 44.07 0.665
Mesh 0.829 24.15 0.121 40.11 0.675
Table 3: Control-signal evaluation on the DexYCB [10] dataset, 256$\times$256 resolution. Using binary masks does not incur significant performance loss compared to joints or meshes.
Occl. SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ Motion Fidel. $\uparrow$
Mask 0.847 24.75 0.121 41.99 0.670
Joints 0.846 24.72 0.122 41.17 0.676
Mesh 0.847 24.83 0.121 42.26 0.665
Mask 0.847 24.79 0.121 41.18 0.672
Joints 0.846 24.69 0.122 41.41 0.676
Mesh 0.848 24.86 0.119 38.83 0.680
Table 4: Control-signal evaluation on the DexYCB [10] dataset, 256$\times$384 resolution. Using binary masks does not incur significant performance loss compared to joints or meshes.
We train 12 different versions of InterDyn on the DexYCB dataset, each version being a combination of the following: resolution (256$\times$256 or 256$\times$384); control signal (mask/joints/mesh); and control signal occlusion (yes/no). In this experiment, as Tab. 3 and Tab. 4 show, we find that the control signal has almost no effect on the image quality, spatiotemporal dynamics, and motion fidelity. This highlights the potential of using simple and easily accessible control signals to generate high-quality video outputs.

### v3

@@ 4.2 Metrics
Image quality. We report the Structural Similarity Index Measure (SSIM[79], Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS[102], Fréchet Inception Distance (FID[30] and unbiased Kernel Inception Distance (KID[4].
Spatio-temporal similarity. To assess the spatio-temporal perceptual similarity between the ground truth and the generated video distributions, we use the Fréchet Video Distance (FVD) and unbiased Kernel Video Distance (KVD) proposed in [74]. We use the implementation of [69].
Motion Fidelity. Through InterDyn, we do not have explicit control over object dynamics, which means that the pixel alignment of an object in the generated and ground truth video is only guaranteed in the starting frame. In this case, comparing generated object motion to the ground truth naively might misrepresent the true quality of object motion over time. Therefore, we adapt the Motion Fidelity (MF) metric proposed by Yatim et al. [93], which measures the similarity between point-tracking trajectories.
To compute the metric for any video, we obtain a mask of the object in the starting frame, sample 100 points on the object, and track these throughout both the ground truth and generated video using CoTracker3 [41]. Given the resulting two sets of tracklets $\mathcal{T}=\{\tau_{1},\dots,\tau_{n}\},\tilde{\mathcal{T}}=\{\tilde{\tau}_{1},\dots,\tilde{\tau}_{m}\}$ the motion fidelity metric is defined as:
@@
with the correlation between two tracklets $\textbf{corr}(\tau,\tilde{\tau})$ [50]:
$\textbf{corr}(\tau,\tilde{\tau})\frac{1}{F}\sum_{k=1}^{F}\frac{{v_{k}^{x}a\cdot\tilde{v}_{k}^{x}+v_{k}^{y}\cdot\tilde{v}_{k}^{y}}}{{\sqrt{(v_{k}^{x})^{2}+(v_{k}^{y})^{2}}\cdot\sqrt{(\tilde{v}_{k}^{x})^{2}+(\tilde{v}_{k}^{y})^{2}}}}\text{,}$ (2)
where $(v_{k}^{x},v_{k}^{y}),(\tilde{v}_{k}^{x},\tilde{v}_{k}^{y})$ are the $k^{th}$ frame displacement of tracklets $\tau,\tilde{\tau}$ respectively. If there are less than 100 points to query on the object due to it being too small, we do not consider the video for the motion fidelity metric.
@@ 4.3 Probing Dynamics with Object Collision Events
Here, we fine-tune InterDyn on an object collision dataset to probe its ability to generate realistic object interactions. Qualitatively, we review whether InterDyn can produce plausible object motion for uncontrolled objects, given the motion of objects entering the scene. In addition, we examine whether InterDyn can generate counterfactual videos for the same input image, but different control signals.
@@
Dataset. We use CLEVRER [97], which provides 20,000 videos of colliding objects with annotated segmentation masks and metadata on collision events. We construct a control signal for objects entering the scene and aim to use InterDyn to generate the motion of the objects that are already present, upon collision. Stationary objects do not receive any form of control signal. Colored masks help the model distinguish unique objects. The frames are cropped and scaled to $320\times 448$, and we only sample input frames before collisions between objects in the scene, to maximize InterDyn’s exposure to interactive dynamics.
@@
Force propagation. InterDyn can generate force propagation dynamics between a controlled object and an uncontrolled object, as well as amongst uncontrolled objects, as illustrated in Fig. 4(a). Here, the red cylinder at the top left is the driving force. It collides with the uncontrolled blue cylinder, which then collides with the uncontrolled purple cylinder, in turn striking the uncontrolled purple cube on the far right. Point-tracking trajectories display how collisions alter each object’s position. This suggests that InterDyn possesses an implicit understanding of physical interactions, enabling it to generate plausible dynamics.
Counterfactual dynamics. By altering the control signal, InterDyn can simulate counterfactual scenarios for the same input image, as shown in Fig. 4(b). In “Generation 1”, the gray cylinder (controlled) collides with the stationary red sphere (uncontrolled), causing it to move; it is later struck by the brown cylinder (controlled), altering its path once again. In “Generation 2”, removing the brown cylinder lets the red sphere continue along its original trajectory, consistent with expectations. Crucially, there is no control signal for the red sphere throughout the sequence; its motion and trajectory are entirely generated by InterDyn.
Probing InterDyn on CLEVRER highlights its ability to generate interactive dynamics for objects within a simple synthetic environment. We provide additional results in video format on our webpage.
Figure 5: Qualitative comparison. A two-state approach such as CosHand [70] struggles with post-interaction object dynamics.
Figure 6: Robustness to noise. SAM2 outputs noisy/coarse masks for frames with considerable motion blur (orange/red). Despite this, InterDyn can generate plausible hand and object dynamics.
[…]
@@ 4.4 Generating Human-Object Interactions
For this experiment, we fine-tune InterDyn on a human-action dataset, focused on hand-object interaction. We encode human movement as a sequence of pixel-aligned binary hand masks and task InterDyn to generate a video with hand movements and corresponding object dynamics.
@@
Dataset. We use Something-Something-v2 (SSV2), which provides 220,847 videos of humans performing basic actions with everyday objects. It contains actions such as “pushing [something] from left to right”, “squeezing [something]” and “lifting [something] with [something] on it”. This dataset allows us to train InterDyn at a larger scale and compare with our baseline CosHand [70]. We train one version of InterDyn at the same resolution as CosHand, 256$\times$256, and a second version at 256$\times$384, which aligns better with the dynamic prior of SVD.
We generate binary hand masks by prompting Segment Anything 2 (SAM2[65] with hand bounding boxes, provided by the Something-Else dataset [59]. Similar to our baseline CosHand, we exclude the “pretending” class. Additionally, we remove all “[something] is done behind [something]” classes since the object would be out of view of the camera. We remove videos smaller than the target resolution/length, ensure the hand and objects are continuously visible, and crop videos larger than the target resolution while centering the object. We include videos without obvious motion and state transition, such as “holding [something]”. Since bounding box annotations are only provided for the train split (79,043 samples for 256$\times$384 and 104,260 for 256$\times$256) and validation split (8,667 samples for 256$\times$384 and 11,229 for 256$\times$256), we report all evaluation metrics on the validation split.
@@
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ FVD $\downarrow$ KVD $\downarrow$ Motion Fidelity $\uparrow$ [93]
Seer [23] 0.418 10.71 0.588 33.35 0.030 287.46 81.31
DynamiCrafter [89] 17.48 0.014 204.11 31.81
CosHand-Independent [70] 0.615 16.87 0.313 2.95 0.002 91.18 19.24 0.432
CosHand-Autoregressive [70] 0.531 14.92 0.408 12.66 0.012 90.30 13.68 0.570
Ours 256$\times$256 0.664 18.60 0.260 4.95 0.004 19.27 1.99 0.633
Ours 256$\times$384 0.680 19.04 0.252 5.07 0.004 22.22 2.09 0.641
Table 1: Quantitative comparison on Something-Something-v2. We compare against two language-instructed video generation methods, Seer [23] and DynamiCrafter [89] and two video extensions of our baseline CosHand [70]. We report results for InterDyn at two resolutions: 256$\times$256 (matching CosHand) and 256$\times$384 (matching SVD’s prior). Methods denoted with $\dagger$ do not use SSV2 as their training dataset.
Baselines. To generate videos using CosHand [70] rather than state transitions, we run CosHand in a frame-by-frame approach (CosHand-Independent) and an auto-regressive approach (CosHand-Autoregressive). In the frame-by-frame variant, each future frame is independently predicted from the initial frame and its corresponding mask:
$\hat{x}_{t+1}=\text{\mbox{CosHand}}(x_{0},h_{0},h_{t+1})\text{,}\quad\forall t\in[0,13],$
where $h_{0}$ and $h_{t+1}$ denote masks, and $x_{0}$ the initial frame. In the auto-regressive variant, we use CosHand to generate video frames sequentially:
$\hat{x}_{t+1}=\text{\mbox{CosHand}}(\hat{x}_{t},h_{t},h_{t+1})\text{,}\quad\forall t\in[0,13],$
with $\hat{x}_{t}$ being the previously generated frame and $\hat{x}_{0}=x_{0}$.
We also compare against two language-instructed video generation methods: Seer [23] and DynamiCrafter [89], which we prompt with a first frame and its corresponding SSV2 class label as instruction. Note that DynamiCrafter was trained with a random video frame as conditioning, i.e. not always the first frame. At inference, its generated videos are thus not strictly a continuation of the conditioning, which precludes frame-aligned metrics. We compare InterDyn against these baselines in Tab. 1.
@@
Quantitative analysis. The frame-by-frame variant of CosHand achieves high image quality but struggles with temporal coherence and motion fidelity. Qualitatively, we notice that object locations are inconsistent across frames. In contrast, the auto-regressive variant improves object motion fidelity but suffers from lower frame-wise image quality due to error propagation. Both variants fail in scenarios requiring accurate post-interaction dynamics, such as when objects continue moving after being released from direct hand contact, as shown in Fig. 5.
Our method, InterDyn (256$\times$384), achieves the best overall performance, surpassing our baseline CosHand in spatio-temporal dynamics, motion fidelity, and all but two image quality metrics. We hypothesize that this might be due to two reasons (1) we use a frozen U-Net, while CosHand fine-tunes its model on SSV2, so CosHand might generate frames closer to the SSV2 distribution, and 2) when SVD was trained, it was initialized with SD weights as spatial layers, and then fine-tuned over multiple stages; this might have degraded its spatial prior, and by extension the quality of produced images compared to CosHand.
Fine-tuned on noisy masks and leveraging its temporal-aware control branch, InterDyn can interpret a noisy control sequence; e.g. when SAM2 produces a coarse and noisy hand mask sequence, InterDyn generates detailed hands including individual fingers, see Fig. 6. Though not always consistent, InterDyn is capable of depicting post-interaction dynamics, such as rolling or sliding objects.
Qualitative analysis. We present diverse qualitative results generated by InterDyn in Fig. 7. Row 1 shows how InterDyn generates the articulated motion of an object. Row 2 showcases pouring water into a glass; note how the water level increases over time. Row 3 demonstrates an object being dropped, moving out of frame when falling, and rolling back in frame once hitting the floor, featuring realistic motion blur synthesis. Rows 4 and 5 illustrate how InterDyn handles squeezing interactions—the rubber and the spring are compressed and restored accordingly. Row 6 demonstrates an understanding of physical size and distance to the camera, as the phone moves closer to the viewer. These results highlight the complexity that InterDyn can handle, implying its generalization ability and future potential as an implicit, yet generalized physical simulator and renderer.
5 Conclusion
We introduced InterDyn, a framework that generates videos of interactive dynamics using large video generation models as implicit physics simulators. By incorporating an interactive control mechanism, InterDyn produces plausible, temporally consistent videos of object interactions—including complex human-object interactions—while generalizing to unseen objects. Our evaluations demonstrate that InterDyn effectively captures continuous motion and subsequent dynamics, outperforming baselines that focus on single future states. This work highlights the potential of using generative video models as physics simulators without explicit reconstruction, opening new avenues for future research.
Acknowledgements: We thank Benjamin Pellkofer for IT support, Peter Kulits for discussions, Tomasz Niewiadomski for coding support, and Allan D. Jepson, Artur Grigorev, Soubhik Sanyal, Omid Taheri & Angjoo Kanazawa for feedback. DT is supported by the ERC Starting Grant (project STRIPES, 101165317).
Disclosure: https://files.is.tue.mpg.de/black/CoI_CVPR_2025.txt DT has received a research gift fund from Google.
Supplementary Material
Control Occlusion SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ Motion Fidelity $\uparrow$ [93]
256$\times$256 256$\times$384 256$\times$256 256$\times$384 256$\times$256 256$\times$384 256$\times$256 256$\times$384 256$\times$256 256$\times$384
Mask 0.829 0.847 24.08 24.75 0.123 0.121 39.94 41.99 0.666 0.670
Joints 0.827 0.846 24.00 24.72 0.124 0.122 40.02 41.17 0.673 0.676
Mesh 0.828 0.847 24.14 24.83 0.122 0.121 41.99 42.26 0.663 0.665
Mask 0.829 0.847 24.15 24.79 0.122 0.121 37.64 41.18 0.675 0.672
Joints 0.827 0.846 24.05 24.69 0.124 0.122 44.07 41.41 0.665 0.676
Mesh 0.829 0.848 24.15 24.86 0.121 0.119 40.11 38.83 0.675 0.680
Table S1: Ablation on control signal. We train and evaluate InterDyn on DexYCB [11]. We ablate: type of control signal (mask, joints, and a colored mesh rendering), presence of object occlusions in the control signal, and two image resolutions (256$\times$256 & 256$\times$384).
Appendix A Ablation study
We use binary hand masks as our control signal due to their widespread availability via methods such as GroundingDINO [52] and SAM2 [65]. However, other control signals—such as skeletons or meshes—might provide richer controllability, since they encode pseudo-3D information and fine-grained correspondences across frames.
Figure S1: Evaluated control signals. From top to bottom: binary mask, joints in the style of OpenPose [9], and colored mesh [61]. Left: w/o object occlusions. Right: w/ object occlusions.
We evaluate the impact of our control signal on performance using the DexYCB dataset [11]. It provides 8,000 videos of hands grasping an object, along with ground-truth 3D hand/object poses/meshes. DexYCB uses the parametric human hand model MANO [67], rendered here as (i) a binary mask (similar to our SSV2 control signal), (ii) joints similar to OpenPose [9], and (iii) a colormap based on [61], see Fig. S1. Additionally, when generating hand masks for SSV2 with SAM2, hand-held objects provide an object contour in the hand mask, inadvertently providing InterDyn with information on the trajectory and shape of the object (a limitation we share with our baseline CosHand [70]). To evaluate its impact, we train separate InterDyn versions on DexYCB, where we render the control signal either with or without the contour of hand-held objects, see Fig. S1.
We present the ablation results in Tab. S1, which indicate that both the type of control signal and contour-leaking effect have minimal impact on image quality, spatio-temporal dynamics, and motion fidelity. These findings softly hint that maintaining hand consistency and driving object interactions does not heavily depend on detailed control signals, rather it does on the video generation model’s implicit understanding of interactive dynamics. This highlights the potential of using simple, readily available control signals to generate high-quality video outputs.
Appendix B State comparison
For completeness, we also compare against CosHand [70] for the second frame of each video and compare these results with the baselines reported in [70], see Tab. S2. For MCVD, UCG, IPix2Pix, TCG, and CosHand (the first five rows), we adopt the results reported in [70] without retraining. Since CosHand does not provide, nor specify, an exact validation split, these numbers are not directly comparable. We also run CosHand on our own validation set.
@@ Appendix B State comparison
Table S2: State comparison on the SSV2 dataset. We compare InterDyn with CosHand and other static baseline methods for generating a single future frame. We report results for InterDyn at two resolutions: 256×256 (matching CosHand) and 256×384 (matching SVD’s prior).